OpenStudy (helpmeee123):

What is the solution to this system of equations? 9x - 2y =15 4x+ 3y= -5

OpenStudy (kohai):

You can use this one of two ways, but I prefer substitution because I find it to be easier

OpenStudy (kohai):

Do you have a preferred method?

OpenStudy (anonymous):

do you know the multiplication method?

OpenStudy (kohai):

There's not really a "multiplication method" -- there's substitution and the addition stuff only

OpenStudy (anonymous):

that's what I meant...my bad

OpenStudy (anonymous):

elimination method

OpenStudy (kohai):

Yeah there you go :) By the way, the best way to master something is to teach it to someone else. Give it a go!

OpenStudy (anonymous):

I will try

OpenStudy (kohai):

I'll stick around to make sure everything goes smoothly :p

OpenStudy (anonymous):

9x - 2y =15 4x + 3y = -5 I like to start by switching the first two variables around (9x and 4x)...like this (4) 9x -2y = 15 (9) 4x + 3y = -5 now multiply all the way across

OpenStudy (anonymous):

(4) 9x -2 =15 36 -8 60 now you try the bottom row

OpenStudy (kohai):

You're trying to get like terms to solve for x, I understand. It will be easier to just simplify as is because y will be by itself. Finding like terms like that will complicate it if you can already solve for y. So you should only do it if you can't just solve for a variable like this

OpenStudy (kohai):

I guess you could eliminate x but it still gets messy...